OpenStudy (anonymous):

cosx-sinx)^2+(cosx+sinx)^2=2 Please help will give medal

OpenStudy (anonymous):

heres what i have so far (cosx - sinx)^2 = cos^2x -2cosxsinx +sin^2x (cosx+sinx)^2 = cos^2x+2cosxsinx+sin^2x

OpenStudy (anonymous):

i need help proving the identity

Nnesha (nnesha):

good. \[\large\rm \color{Red}{(cosx-sinx)^2}+\color{blue}{(cosx+sinx)^2}\] \[\large\rm \color{Red}{ \cos^2x -2cosxsinx +\sin^2x} +\color{blue}{ \cos^2x+2cosxsinx+\sin^2x}\] there is plus sign(addition ) so we don't need to put parentheses now just combine like terms

Nnesha (nnesha):

a*

OpenStudy (anonymous):

cosxsinex

Nnesha (nnesha):

hmm what about them ??? what are like terms in that equation ?

OpenStudy (anonymous):

the like terms would be cosine

Nnesha (nnesha):

it's not cos cos^2x and what else ?

OpenStudy (anonymous):

sin^2x

Nnesha (nnesha):

alright what else ?

OpenStudy (anonymous):

=1

Nnesha (nnesha):

hmm i'm saying what are like terms in this equation ? . \[\large\rm \color{Red}{ \cos^2x -2cosxsinx +\sin^2x} +\color{blue}{ \cos^2x+2cosxsinx+\sin^2x}\]

Nnesha (nnesha):

do you know what are `like terms ` ?

OpenStudy (anonymous):

like terms are equations that have the same varible

Nnesha (nnesha):

correct! combine like terms in that equation

OpenStudy (anonymous):

1-2cosxsinx+cos^2x+sin2x

Nnesha (nnesha):

\[\large\rm \color{Red}{ \cos^2x} -2cosxsinx +\sin^2x +\color{red}{ \cos^2x}+2cosxsinx+\sin^2x\] cos^2(x) and cos^2(x) are like terms so combine them cos^2(x)+cos^2(x)= ???

OpenStudy (anonymous):

cos^4x

Nnesha (nnesha):

no lets review `like terms` how would you combine \[\large\rm x+x = ??\]

OpenStudy (anonymous):

i would xx+xx

OpenStudy (anonymous):

xx+xx

Nnesha (nnesha):

please explain what is that ? how did you get 2 x on each side of plus sign

OpenStudy (anonymous):

i went by your example

OpenStudy (anonymous):

it would just be cos

Nnesha (nnesha):

when we combine `like terms` we should add/subtract their coefficients\[\large\rm 1x+1x= 2x\] when we `multiply like terms` then we should add their exponents

OpenStudy (anonymous):

2cos

Nnesha (nnesha):

right but we should keep the square \[\huge\rm \cos^2x+\cos^2x= 2\cos^2x\] \[\large\rm \color{Red}{ 2\cos^2x} -2cosxsinx +\sin^2x \color{red}{}+2cosxsinx+\sin^2x\] now combine other like terms

OpenStudy (anonymous):

cosx+2sinx

Nnesha (nnesha):

how did you get cos x ??

OpenStudy (anonymous):

i added cosx-cosx sin+sinxcosxsinx+sinx

Nnesha (nnesha):

here are some example \[\large\rm \rm 2x^2+4x^2\] the variables of both terms are the same and same exponent so i can combine their `coefficients` \[(2+4)x^2=7x^2\] i just added coefficients and keep the x^2 variable

Nnesha (nnesha):

let cos x = a and sin x = b alright now i'm gonna replace alll sin x with b and cosx with a

Nnesha (nnesha):

\[\large\rm \color{red}{ a^2 -2ab+b^2} +\color{blue}{ a^2+2ab+b^2}\] try to combine like terms now

OpenStudy (anonymous):

cosx-cosinxsin+sinx+cosx+cosxsin+sinx

Nnesha (nnesha):

\[\large\rm \color{red}{ a^2 -2ab+b^2} +\color{blue}{ a^2+2ab+b^2}\] what are like terms in this equation ?

OpenStudy (anonymous):

cos sin

Nnesha (nnesha):

i replaced cosx with a and sin x = b to make it looks better and simple

Nnesha (nnesha):

\[\large\rm \color{red}{ a^2 -2ab+b^2} +\color{blue}{ a^2+2ab+b^2}\] so what are like terms in this equation

Nnesha (nnesha):

remember x and x^2 are NOT like terms

OpenStudy (anonymous):

cos2 and sin2???

Nnesha (nnesha):

alright then keep sin and cos \[\large\rm \color{Red}{ \cos^2x} \color{pink}{-2cosxsinx }+\color{blue}{\sin^2x} +\color{red}{ \cos^2x}+\color{pink}{2cosxsinx}+\color{blue}{\sin^2x}\] these are like terms

Nnesha (nnesha):

sin^2x+sin^2x= ?? cos^2x+cos^2= ?? 2sinxcosx - 2sincosx= ??

OpenStudy (anonymous):

2sin^2x 2cos^2x sin2x−2sinxcosx

OpenStudy (anonymous):

am i right

Nnesha (nnesha):

first two are right 2sinxcos-2sinxcos is same as 2ab -2ab = ??

OpenStudy (anonymous):

sin2x−2cosxsinx

Nnesha (nnesha):

now what happen when we subtract like terms like a-a= x-x= ? 2-2= ?

OpenStudy (anonymous):

we get 0

Nnesha (nnesha):

right so \[\rm 2sinxcosx - 2sinxcosx =?? \] you're subtracting like terms

OpenStudy (anonymous):

sin2x-2sinxcosx ?

Nnesha (nnesha):

how did you get that ?

OpenStudy (anonymous):

i subtracted

Nnesha (nnesha):

we know cos^2+cos^2 = 2cos ^2 \[\large\rm \color{Red}{ 2\cos^2x} -2cosxsinx +\sin^2x \color{red}{}+2cosxsinx+\sin^2x\] and sin^2x+sin^2x= 2sin^2x \[\large\rm \color{Red}{ 2\cos^2x} -2cosxsinx +\color{blue}{2sin^2x}\color{red}{}+2cosxsinx\] now we need to combine 2sinxcosx-2sinxcosx

OpenStudy (anonymous):

Would it be 0

Nnesha (nnesha):

right so we left with \[\huge\rm 2\cos^2x+2\sin^2x\] now take out the common factor

OpenStudy (anonymous):

1

Nnesha (nnesha):

what about one ??

OpenStudy (anonymous):

That's what it would equal

Nnesha (nnesha):

hmm now first we take out the common factor

Nnesha (nnesha):

no** not now

Nnesha (nnesha):

\[\huge\rm 2\cos^2x+2\sin^2x\] 2 is common so we should take it out

OpenStudy (anonymous):

So I would get cosx+sinx

Nnesha (nnesha):

no

Nnesha (nnesha):

cos^2x and sin^2 not just cos sin that's very big mistake

Nnesha (nnesha):

\[\huge\rm 2(\cos^2x+\sin^2x)\]and we should keep the common factor outside the parentheses

Nnesha (nnesha):

now use the special identity

OpenStudy (anonymous):

I would get 2

Nnesha (nnesha):

right

OpenStudy (anonymous):

So would my identity be false??

Nnesha (nnesha):

why ?

OpenStudy (anonymous):

Or true

Nnesha (nnesha):

what do you think ? we were working on left side is that equal to right side ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Wait no

Nnesha (nnesha):

no to what ?

OpenStudy (anonymous):

It's not an identity

Nnesha (nnesha):

answer my question is the left side equal to right side ?

OpenStudy (anonymous):

Yes

Nnesha (nnesha):

so if both sides are equal then the equation is correct true identity if both sides aren't equal like `4=6` then false

OpenStudy (anonymous):

Ok I got thank you again

Nnesha (nnesha):

np :=))